Calculate the value of the translational partition function of O2 at 1000 K for a volume of 1 m3. Submit your answer multiplied by 10−32 (i.e. if you calculate 4.2e33, input 42).
1
Expert's answer
2020-06-11T11:41:01-0400

Here we have to find the partition function of oxygen we know that mass of oxygen is = 32

Now we will calculate the de borglies thermal wavelength using formula


Λ="\frac{h}{\sqrt {2 \pi m k_B T}}"

Here, T= 1000 K, h= plancks constant, "k_B=" Boltzmann constant


Λ="\frac{6.67\times 10 ^{-34}}{\sqrt {2 \times 3.14 \times 2\times32\times 1.67 \times 10^{-27} \times 1.38 \times 10^{-23}\times 1000}}"


Λ= "6.9 \times 10 ^{-12}"

And for the partition function we know that

"q_{tr}= \frac{V}{Λ^3}= \frac{1}{(6.9\times 10 ^{-12})^3} = 30.38 \times 10 ^{32}" = so here we will input only 30 as per the question